Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company has a standard cost system in use for all of its products. According to the standards that have been set for the seat covers, the factory should work 1,015 hours each month to produce 2,030 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 59,276 $ 29.20
Direct labor $ 8,120 4.00
Variable manufacturing overhead (based on direct labor-hours) $ 3,857 1.90
$ 35.10

During August, the factory worked only 700 direct labor-hours and produced 1,500 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (8,400 yards) $ 42,000 $ 28.00
Direct labor $ 6,300 4.20
Variable manufacturing overhead $ 3,150 2.10
$ 34.30

At standard, each set of covers should require 4.0 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

This is the last question in the assignment. To submit, use Alt + S. To access other questions, proceed to the question map button.Next
Answer #1

1) Material price variance = (7.3*8400-42000) = 19320 F

Material quantity variance = (1500*4-8400)*7.3 = 17520 U

2) Labor rate variance = (8*700-6300) = 700 U

Labor efficiency variance = (750-700)*8 = 400 F

3) Variable overhead rate variance = (3.8*700-3150) = 490 U

Variable overhead efficiency variance = (750-700)*3.8 = 190 F